OpenStudy (anonymous):

What is the value of the variable that makes the statement true? a + (–7) + 3 = 0

OpenStudy (anonymous):

a + (-7) + 3 = 0 a - 7 + 3 = 0 a -4 = 0 a = 4 Thus, 4 is the value of the variable which satisfies the equation.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

@Studygirl97

OpenStudy (anonymous):

@Teleprto101 You are welcomed.